OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

As it is parallel to y = x + 9 it will have the same slope, also 1, so you substitute -2 and -3 in the general equation for a line and you get: -3 = -2 + b, so b = -1 Your line is y = x - 1

OpenStudy (anonymous):

standard form is y-x=-1

OpenStudy (anonymous):

x – y = 1 x + y = 1 x – y = –5 x + y = –5 are the options

OpenStudy (anonymous):

y=x-1 is slope intercept form